OpenStudy (anonymous):

Show Steps: y – 31 = -5/24(x – 15) Solve for y

OpenStudy (hba):

@SephI Show me the working and i tell you where you go wrong...........

OpenStudy (anonymous):

y - 31 = -5/24(X - 15) y - 31 = -5/24x + (Don't know how to do this part)

OpenStudy (hba):



OpenStudy (hba):

@SephI This Is an example ^

OpenStudy (hba):

Now try to solve your question @SephI

OpenStudy (anonymous):

-5/24 * 15 <-- I dunno how

OpenStudy (hba):